OpenStudy (anonymous):

Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. z1 = 1/8 (cos 2pi/3 + i sin 2pi/3) z2= 1/3 (cos pi/4 + i sin pi/4)

ganeshie8 (ganeshie8):

\(\large e^{i\theta} = \cos \theta + i \sin \theta\)

ganeshie8 (ganeshie8):

cos 2pi/3 + i sin 2pi/3 = \(\large e^{i\frac{2\pi}{3}}\)

ganeshie8 (ganeshie8):

convert both complex numbers to exponent form.. and do the usual simplifications..

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (\cos 2\pi /3 + i \sin 2 \pi /3)}{ 1/3 (\cos \pi /4 + i \sin \pi /4) }} \) \(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (e^{ i2 \pi /3})}{ 1/3 (e^{i \pi /4}) }} \) \(\large \mathbb{\frac{z1}{z2} = 3/8 (e^{ i2 \pi /3 - i\pi/4})} \)

ganeshie8 (ganeshie8):

still here ?

OpenStudy (anonymous):

I'm here. Just processing lol

ganeshie8 (ganeshie8):

good, 3rd step above is just simple division of a fraction...

OpenStudy (anonymous):

Sorry I'm super tired (acute insomniac)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (\cos 2\pi /3 + i \sin 2 \pi /3)}{ 1/3 (\cos \pi /4 + i \sin \pi /4) }} \) \(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (e^{ i2 \pi /3})}{ 1/3 (e^{i \pi /4}) }} \) \(\large \mathbb{\frac{z1}{z2} = \frac{3}{8}(e^{ i2 \pi /3 - i\pi/4})} \) \(\large \mathbb{\frac{z1}{z2} = \frac{3}{8} (e^{ i 5 \pi /12})} \)

ganeshie8 (ganeshie8):

change it back to polar form again

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (\cos 2\pi /3 + i \sin 2 \pi /3)}{ 1/3 (\cos \pi /4 + i \sin \pi /4) }} \) \(\large \mathbb{\frac{z1}{z2} = \frac{ 1/8 (e^{ i2 \pi /3})}{ 1/3 (e^{i \pi /4}) }} \) \(\large \mathbb{\frac{z1}{z2} = \frac{3}{8}(e^{ i2 \pi /3 - i\pi/4})} \) \(\large \mathbb{\frac{z1}{z2} = \frac{3}{8} (e^{ i 5 \pi /12})} \) \(\large \mathbb{\frac{z1}{z2} = \frac{3}{8} (\cos 5 \pi /12 + i \sin 5 \pi /12) } \)

ganeshie8 (ganeshie8):

see if that makes some sense..

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

It's a little more clear now :)

ganeshie8 (ganeshie8):

np, glad to hear that :)